# Dynamical probe of the pseudo Jahn-Teller effect in one-dimensional confined fermions

### v1

@@ IV.2.1 Density analysis
The NVABO approximation does not capture this behavior, but rather it exhibits similar dynamics to the previously analyzed cases, see Fig. 10(a2) and 10(b2). Qualitatively the behavior is almost identical to the $g=2.5~{}\hbar\omega_{B}\ell_{B}$ case of Fig. 9(a2) and 9(b2). The main differences are a longer oscillation period and the emergence of more pronounced peak density fluctuations on the impurity species, see Fig. 10(b2). On the other hand the VABO approximation, is able to capture some of the main features of the exact impurity dynamics, compare Fig. 10(a3) and 10(b3) to Fig. 10(a1) and 10(b1). This leads to the conclusion that the trapping of the impurity density outside its bath, that can be seen in Fig. 10(b3), is substantially influenced by the inclusion of the Born-Huang correction, which accounts for changes in the bath’s kinetic energy. Only a small fraction of the impurity density passes and transmits through the bath species, where it shows tunneling-like dynamics. Nevertheless, the majority of the impurity density remains outside the bath and does not undergo well-pronounced oscillations. As a result, in the bath one-body density (shown in Fig. 10(a3)), we observe only a slight density valley during the first half of the initial faint-dipole oscillation. This valley subsequently vanishes, leaving behind only minor perturbations (small wiggles) in the bath density. A collective dipole oscillation, as observed in the numerically exact results, see Fig. 10(a1) is absent for $\rho^{(1)}_{B}(x;t)$.
The above already indicates that in this regime the dynamics of the system tends towards being completely diabatic since the impurity is able to excite a collective mode of the bath, which of course it is difficult to accurately describe with a small number of PEC, let alone a single one as in the adiabatic approximation. Nevertheless, comparing to VABO we can verify that the origin of the localization of the impurity outside its fermionic environment originates from the bath’s kinetic energy contribution, as accounted for by the Born-Huang term.
@@ IV.2.2 Analysis in terms of PEC
Building on our earlier analysis of smaller shifts, we once again examine the populations of the PECs to deepen our understanding of the one-body density trends shown in Fig. 2, 9, and 10. This allows us to further unravel the role of beyond Born-Oppenheimer effects, with particular emphasis on the PJTE. To achieve this, we analyze the contributions of individual PECs via Eq. (12), using Fig. 11 to illustrate their behavior and its implications. In the simple non-interacting case, as shown in Fig. 11(a), we observe, as expected, that only the lowest PEC is involved. This corroborates our findings on the one-body density (Fig. 2(b2)), confirming that this behavior is entirely described within the adiabatic approximation.
@@
To proceed further, we examine the impurity density contributions of the different PECs in Fig. 12. Of course, for $g=0$ our results are straightforward, since only the $j=1$ PEC is involved and thus $\rho^{(1)}_{I,1}(x;t)=\rho^{(1)}_{I}(x;t)$, see Fig. 12(a1) and Fig. 12(b1). In the weakly interacting case, we observe a qualitatively similar behavior to the case of smaller displacements. Indeed, $\rho^{(1)}_{I,1}(x;t)$ exhibits a structure similar to the impurity density within VABO, as seen in Fig. 12(a2) to  9(b3), however, the amplitude of $\rho^{(1)}_{I,1}(x;t)$ is significantly reduced due to the property $n_{1}(t)=\int{\rm d}x~{}\rho^{(1)}_{I,1}(x;t)$ of the PEC densities and the drastic reduction of $n_{1}(t)$ during the dynamics, see Fig. 11(b). Turning to the case of the excited PEC we observe that the case of $x_{s}=1.5\ell_{B}$ shows a similar qualitative behavior as the case of $x_{s}=0.5\ell_{B}$, compare Fig. 12(b2) to Fig. 6(b2). In particular, a density component of the second PEC emerges near each avoided crossing as the impurity density of the ground PEC goes through this crossing, which indicates diabatic transport. Similarly also to the case of $x_{s}=0.5\ell_{B}$ this impurity density fraction exhibits tunneling dynamics from one crossing to another, an effect that is more prominently visible for $t\approx 6.5\omega_{B}^{-1}$ and $\ell_{B}\leq x\leq 2l_{B}$. The above results suggest that despite the fact that in this case the dynamics of the system deviates substantially from the adiabatic case, via the generation of collective excitations observable in the bath density, the underlying cause is the diabatic transport of the impurity density. This shows that the non-adiabatic behavior of the system observed for both $x_{s}=0.5\ell_{B}$ and $x_{s}=1.5\ell_{B}$ in the $g=2.5\hbar\omega_{B}\ell_{B}$ case essentially defines a crossover region from fully adiabatic behavior for very low interactions and the collective excitation behavior for strong interaction, which we examine below.
@@
To finish this section we briefly discuss the occupation of different vibrational states of the two lowest PEC. In the non-interacting case, the shift in potential energy leads to the occupation of higher-lying states of the ground-state PEC at the center of the trapping potential (Fig. 13(a)). With the increased displacement, $x_{s}=1.5$ we observe that the first excited state of the ground-state PEC initially becomes the most populated as the Poisson distribution of the eigenstate population within the initial coherent state becomes more apparent. Due to the absence of interactions, the bath and impurity remain separable, resulting in no temporal evolution of the state populations. Upon introducing weak interactions ($g=2.5\hbar\omega_{B}\ell_{B}$), as depicted in Fig. 13(b) and (d), we initially observe a behavior in the ground-state PEC similar to the non-interacting case but shifted to higher lying eigenstates, owing to the extra-shift of the initial impurity configuration due to the interaction. However, once the system encounters an avoided crossing—evidenced by the still-discernible dipole oscillation in the impurity’s one-body density (Fig. 9(b1))—the second to fifth eigenstates of the first excited PEC become significantly populated. Notably, in contrast to the smaller displacement case, these states (including the third state) remain occupied over time. The effective potential of the first excited PEC, when combined with the structure of the corresponding eigenfunctions, reveals that the triple-well structure facilitates trapping, whereas the weakly pronounced double-well structure in the ground-state PEC does not. Moving to the strongly interacting regime ($g=5.0\hbar\omega_{B}\ell_{B}$), we observe a similar pattern. In particular, during time-instances where the impurity density penetrates the trap center despite the repulsive interaction, the first excited state in the excited PEC dominates, as shown in Fig. 13(c). In contrast, the lowest state in the ground-state PEC remains effectively unoccupied. This behavior correlates with the oscillation of the impurity’s one-body density, which extends beyond $x>2.0\ell_{B}$, with only a small fraction reaching the trap center. Examining the ground-state PEC’s zeroth eigenfunction (Fig. 7(b3)), we note its localization at the trap center, yet the strong repulsion prevents significant impurity presence in this region. Finally, a key point of Fig. 7(a2), 7(b2), 7(a3), and 7(b3) is that in the interacting case the occupation of states in the excited PEC lie in the same energy range as the lowest lying PEC.
@@ V Stronger trapping confinement of the impurity
As part of our ground-state analysis, we explored the dynamics of a tightly confined impurity species in depth [24]. Building on this foundation, we now examine the dynamic behavior for this scenario $\omega_{I}=4\omega_{B}$, focusing on the case where the avoided crossing at $x=0$ occurring between the lowest and first excited PECs becomes not visible because of the large contribution of the impurity confining potential, compare Fig. 14(b) to 14(a).
@@ VI Summary and Outlook
These results provide significant insights into non-adiabatic processes in ultracold quantum gases and establish a theoretical foundation for experimental investigations in controlled atomic settings. The ability to tune interactions via Fano-Feshbach resonances or confinement-induced resonances opens new possibilities for experimentally realizing quantum-controlled impurity dynamics.
@@
Finally, our findings naturally connect to broader concepts in many-body physics. In particular, a fascinating question is the potential link between the non-adiabaticity observed here and the well-known Anderson orthogonality catastrophe [52, 53]. Investigating how impurity-induced decoherence and correlation effects manifest dynamically in strongly interacting fermionic systems could provide new insights into quantum impurity problems, polaron physics, and emergent collective behavior in ultracold gases.

### v2

@@ IV.2.1 Density analysis
The NVABO approximation does not capture this behavior, but rather it exhibits similar dynamics to the previously analyzed cases, see Fig. 11(a2) and 11(b2). Qualitatively the behavior is almost identical to the $g=2.5~{}\hbar\omega_{B}\ell_{B}$ case of Fig. 10(a2) and 10(b2). The main differences are a longer oscillation period and the emergence of more pronounced peak density fluctuations on the impurity species, see Fig. 11(b2). On the other hand, the VABO approximation, is able to capture some of the main features of the exact impurity dynamics, compare Fig. 11(a3) and 11(b3) to Fig. 11(a1) and 11(b1). This leads to the conclusion that the trapping of the impurity density outside its bath, that can be seen in Fig. 11(b3), is substantially influenced by the inclusion of the Born-Huang correction, which accounts for changes in the bath’s kinetic energy. Only a small fraction of the impurity density passes and transmits through the bath species, where it shows tunneling-like dynamics. Nevertheless, the majority of the impurity density remains outside the bath and does not undergo well-pronounced oscillations. As a result, in the bath one-body density (shown in Fig. 11(a3)), we observe only a slight density valley during the first half of the initial faint-dipole oscillation. This valley subsequently vanishes, leaving behind only minor perturbations (small wiggles) in the bath density. A collective dipole oscillation, as observed in the numerically exact results, see Fig. 11(a1) is absent for $\rho^{(1)}_{B}(x;t)$.
The above already indicates that in this regime the dynamics of the system tends towards being completely diabatic since the impurity is able to excite a collective mode of the bath, which of course it is difficult to accurately describe with a small number of PEC, let alone a single one as in the adiabatic approximation. Nevertheless, comparing to VABO we can verify that the origin of the localization of the impurity outside its fermionic environment originates from the bath’s kinetic energy contribution, as accounted for by the Born-Huang term. Fig. 7(b) offers a graphical illustration of this interplay, showing the qualitative behavior of both the Born-Huang term, which is shifted by a constant factor, and the PECs. It provides an intuitive summary of the spatial dynamics of the impurity, as revealed by the various approaches discussed earlier. The alignment between the positions of the avoided crossings and the peaks of the Born-Huang term $V_{\rm ren}^{11}(x_{I})$ indicates the emergence of an effective harmonic oscillator potential, which leads to the self-trapping of the impurity in the outer region, exhibiting quasi-diabatic behavior dictated by the structure of the PECs. As a result, only a small fraction of the impurity density, as previously observed, is able to escape this trapping.
@@ IV.2.2 Analysis in terms of PEC
Building on our earlier analysis of smaller shifts, we once again examine the populations of the PECs to deepen our understanding of the one-body density trends shown in Fig. 2, 10, and 11. This allows us to further unravel the role of beyond Born-Oppenheimer effects, with particular emphasis on the PJTE. To achieve this, we analyze the contributions of individual PECs via Eq. (12), using Fig. 12 to illustrate their behavior and its implications. In the simple non-interacting case, as shown in Fig. 12(a), we observe, as expected, that only the lowest PEC is involved. This corroborates our findings on the one-body density (Fig. 2(b2)), confirming that this behavior is entirely described within the adiabatic approximation.
@@
To proceed further, we examine the impurity density contributions of the different PECs in Fig. 13. Of course, for $g=0$ our results are straightforward, since only the $j=1$ PEC is involved and thus $\rho^{(1)}_{I,1}(x;t)=\rho^{(1)}_{I}(x;t)$, see Fig. 13(a1) and Fig. 13(b1). In the weakly interacting case, we observe a qualitatively similar behavior to the case of smaller displacements. Indeed, $\rho^{(1)}_{I,1}(x;t)$ exhibits a structure similar to the impurity density within VABO, as seen in Fig. 13(a2) to  10(b3), however, the amplitude of $\rho^{(1)}_{I,1}(x;t)$ is significantly reduced due to the property $n_{1}(t)=\int{\rm d}x~{}\rho^{(1)}_{I,1}(x;t)$ of the PEC densities and the drastic reduction of $n_{1}(t)$ during the dynamics, see Fig. 12(b). Turning to the case of the excited PEC we observe that the case of $x_{s}=1.5\ell_{B}$ shows a similar qualitative behavior as the case of $x_{s}=0.5\ell_{B}$, compare Fig. 13(b2) to Fig. 6(b2). In particular, a density component of the second PEC emerges near each avoided crossing as the impurity density of the ground PEC goes through this crossing, which indicates diabatic transport. Similarly also to the case of $x_{s}=0.5\ell_{B}$ this impurity density fraction exhibits tunneling dynamics from one crossing to another, an effect that is more prominently visible for $t\approx 6.5\omega_{B}^{-1}$ and $\ell_{B}\leq x\leq 2l_{B}$. The above results suggest that despite the fact that in this case the dynamics of the system deviates substantially from the adiabatic case, via the generation of collective excitations observable in the bath density, the underlying cause is the diabatic transport of the impurity density. This shows that the non-adiabatic behavior of the system observed for both $x_{s}=0.5\ell_{B}$ and $x_{s}=1.5\ell_{B}$ in the $g=2.5\hbar\omega_{B}\ell_{B}$ case essentially defines a crossover region from fully adiabatic behavior for very low interactions and the collective excitation behavior for strong interaction, which we examine below.
@@
To finish this section we briefly discuss the occupation of different vibrational states of the two lowest PEC. In the non-interacting case, the shift in potential energy leads to the occupation of higher-lying states of the ground-state PEC at the center of the trapping potential (Fig. 14(a)). With the increased displacement, $x_{s}=1.5$ we observe that the first excited state of the ground-state PEC initially becomes the most populated as the Poisson distribution of the eigenstate population within the initial coherent state becomes more apparent. Due to the absence of interactions, the bath and impurity remain separable, resulting in no temporal evolution of the state populations. Upon introducing weak interactions ($g=2.5\hbar\omega_{B}\ell_{B}$), as depicted in Fig. 14(b) and (d), we initially observe a behavior in the ground-state PEC similar to the non-interacting case but shifted to higher lying eigenstates, owing to the extra-shift of the initial impurity configuration due to the interaction. However, once the system encounters an avoided crossing—evidenced by the still-discernible dipole oscillation in the impurity’s one-body density (Fig. 10(b1))—the second to fifth eigenstates of the first excited PEC become significantly populated. Notably, in contrast to the smaller displacement case, these states (including the third state) remain occupied over time. The effective potential of the first excited PEC, when combined with the structure of the corresponding eigenfunctions, reveals that the triple-well structure facilitates trapping, whereas the weakly pronounced double-well structure in the ground-state PEC does not. Moving to the strongly interacting regime ($g=5.0\hbar\omega_{B}\ell_{B}$), we observe a similar pattern. In particular, during time-instances where the impurity density penetrates the trap center despite the repulsive interaction, the first excited state in the excited PEC dominates, as shown in Fig. 14(c). In contrast, the lowest state in the ground-state PEC remains effectively unoccupied. This behavior correlates with the oscillation of the impurity’s one-body density, which extends beyond $x>2.0\ell_{B}$, with only a small fraction reaching the trap center. Examining the ground-state PEC’s zeroth eigenfunction (Fig. 8(b3)), we note its localization at the trap center, yet the strong repulsion prevents significant impurity presence in this region. Finally, a key point of Fig. 8(a2), 8(b2), 8(a3), and 8(b3) is that in the interacting case the occupation of states in the excited PEC lie in the same energy range as the lowest lying PEC.
@@ V Stronger trapping confinement of the impurity
As part of our ground-state analysis, we explored the dynamics of a tightly confined impurity species in depth [24]. Building on this foundation, we now examine the dynamic behavior for this scenario $\omega_{I}=4\omega_{B}$, focusing on the case where the avoided crossing at $x=0$ occurring between the lowest and first excited PECs becomes not visible because of the large contribution of the impurity confining potential, compare Fig. 15(b) to 15(a).
@@ VI Summary and Outlook
These results provide significant insights into non-adiabatic processes in ultracold quantum gases and establish a theoretical foundation for experimental investigations in controlled atomic settings. The ability to tune interactions via Fano-Feshbach resonances or confinement-induced resonances opens new possibilities for experimentally realizing quantum-controlled impurity dynamics. In particular, in situ imaging provides access to both the impurity and bath position, allowing for a direct comparison to the spatial resolved densities in this work.
@@
Finally, our findings naturally connect to broader concepts in many-body physics. In particular, a fascinating question is the potential link between the non-adiabaticity observed here and the well-known Anderson orthogonality catastrophe [52, 53]. Traditionally, the Anderson orthogonality catastrophe arises when a localized and strongly interacting impurity modifies the many-body wavefunction of its surrounding bath, leading to a vanishing overlap between the interacting and non-interacting ground states in the thermodynamic limit, that is $Z=\sqrt{|\langle\psi_{g=0}|\psi_{g\neq 0}\rangle|^{2}}\propto N^{-\alpha}$ with $\alpha>0$. In our framework, this orthogonality is inherently present from the outset: the multi-channel Born-Oppenheimer ansatz explicitly accounts for the dependence of each bath fermion’s state on the instantaneous position of the impurity. These couplings induce collective excitations and lead to spontaneous symmetry breaking and impurity self-trapping. This constitutes a qualitatively distinct regime of impurity physics, where the breakdown of adiabaticity gives rise to emergent, non-perturbative dynamics beyond the conventional polaron picture, in which the impurity remains delocalized and is dressed by virtual particle hole excitations [54]. In contrast, the pseudo-Jahn-Teller regime features strong, non-adiabatic coupling that results in localized impurity behavior and symmetry-broken configurations. Our results thus provide a foundation for future studies on quasiparticle breakdown and many-body entanglement arising from non-adiabatic impurity motion, offering new insights into quantum impurity problems in ultra-cold atomic systems and their collective behavior.